TOOK TO LONG DONT SOLVE FIGURED IT OUT MYSELF

If you heat a 5.0 liter steel tank to 150.0 degree Fahrenheit that contains 101.2 grams of propane MW 44 grams /mole the pressure in the tank will be?

A. 1300 torr

B. 9.89 atm

C. 9.7 x 10 3 mm of Hg

D. 195 lb/in 2
Answer #1

T= 150 °F (F-32 ) * (ISO - 32 ). . . T= 65.55°C T = (65.55 + 273.15)k T= 338.55k/ moles of propane = 2.3 mol mol = 101.2 fm 4